Find the Laplace transforms of the following function:

  1. L{t^2sinkt}
1
Expert's answer
2022-05-22T17:41:12-0400

We have

"L(sin kt)=\frac{k}{s^2+k^2}"

"L(t^2sin(kt))=(-1)^2\frac{d^2}{s^2}(\frac{k}{k^2+s^2})=\frac{d}{ds}\frac{-2ks}{(k^2+s^2)^2}=\frac{-2k(k^2+s^2)^2-8ks^2(k^2+s^2)}{(k^2+s^2)^4}=\frac{6ks^2-2k^3}{(s^2+k^2)^3}"